A spherical surface completely surrounds a collection of charges. Find the electric flux (with its sign) through the surface if the collection consists of (a) a single +7.20 x 10 C charge, (b) a single -8.20 x 10-6 c charge, and (c) both of the charges in (a) and (b). (a) NumberT8.13E5 (b) Number. (c) Number Click if you Units-TN-rm^ 2/C 수 Units TN-m22/C 수 STN-m^2/C 수 :Unit -1.1E5 is+/-2% ork for this question: Open Show Work
Answer #1

eo = 8.85*10^-12

Flux = q / eo

a) Flux = (+7.20 x 10-6) / eo

= 8.14*10^5 Nm^2/C

b) Flux = (-8.20 x 10-6) / eo

= -9.27E5 Nm^2/C

c) Flux = (7.2 x 10-6) + (-8.2 x 10-6) / eo

= -1.13E5 Nm^2/C